1. calculate Kp for the reaction C(s) + Co2(g) = 2Co(g at 1300 k and 1 atm, the equilibrium mixture contains 85% co and 10.5% co2 by volume.
2. calculate the partial pressure of Co2 if Pco is changed to 10^-3 ATM.
Answer #1

Q1:

C(s) + CO2(g) 2CO(g)

PCO = (85/100)×1atm = 0.85atm

PCO2 = 15/100 × 1atm = 0.15atm

Kp = (PCO)2/PCO2 = (0.85)2/0.15 = 4.817

= 4.82

Q2:

PCO2 =( PCO)2/Kp = (10-3)2/4.817 = 2.076×10-7

= 2.1×10-7atm